# TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2

### 03 Dec 2023 v2

@@ 5.2 In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
Table 4: Comparative analysis of the Efficiency-Accuracy-Complexity Score (EACS) across various anomaly detection models, highlighting the performance of TransNAS-TSAD. The EACS is calculated by normalizing and weighting F1 scores (0.4), training times (0.4), and parameter counts (0.2), demonstrating the balanced efficiency and accuracy of TransNAS-TSAD against benchmark models across multiple datasets
Arch and Training Params SMAP UCR MBA SWaT MSL SMD NAB WADI
Learning Rate 0.0002128 0.0019997 0.003441 0.0000403 0.006192 0.0002273 0.006924 0.0000131
Dropout Rate 0.2353 0.4474 0.1795 0.3836 0.3730 0.1554 0.4560 0.2237
Dim Feedforward 101 124 41 42 121 86 10 16
Batch Size 32 48 16 32 32 48 96 112
Encoder Layers 2 1 3 2 2 3 1 2
Decoder Layers 1 2 1 3 3 1 2 1
Activation Func. sigmoid leaky_relu tanh tanh leaky_relu sigmoid relu sigmoid
Time Warping False True True True True False False False
Time Masking True True False True True False True False
Gaussian Noise 0.000151 0.027956 0.007925 0.058019 0.000428 0.000628 0.000119 0.000194
Linear Embedding False True False False True True True True
Phase Type 2phase iterative iterative 2phase 2phase 2phase iterative 2Phase
Self Conditioning False True False True True False True True
Layer Norm False False False True True True False True
Pos. Enc. Type sinusoidal fourier fourier sinusoidal fourier sinusoidal sinusoidal fourier
FFN Layers 1 1 1 3 2 3 1 3
Attn Heads 25 1 2 51 55 38 1 127
Window Size 10 20 14 22 26 12 18 26
Table 5: TransNAS-TSAD Hyperparameter Optimization Results: Detailed summary of the optimal training and architectural hyperparameters identified for each dataset, demonstrating the adaptive precision of TransNAS-TSAD in time series anomaly detection.
5.2.1 Optimal Model Configurations Identified by TransNAS-TSAD
@@
Table 5 showcases the culmination of TransNAS-TSAD’s NSGA-II optimization process, highlighting the training and architectural hyperparameters of models that achieved the best balance between high F1 scores and low parameter counts. These configurations represent the pinnacle of the framework’s capability to identify models uniquely suited to each dataset’s anomaly detection needs.
@@
A key observation from the table is the diversity in architectural parameters across datasets, reflecting TransNAS-TSAD’s adeptness in tailoring models to specific data characteristics. For example, in the SMAP dataset, the model employs a relatively simple architecture with fewer encoder and decoder layers, which is effective in handling the environmental time series data of this dataset. In contrast, the WADI dataset, known for its complex sensor network, necessitates a more intricate model structure, evident in its higher number of attention heads and the use of a two-phase reconstruction and refinement strategy.
@@
These results from TransNAS-TSAD’s NAS process validate the framework’s ability to not only fine-tune hyperparameters for operational efficiency but also to intricately adapt its architectural design to the nuanced demands of various time series anomaly detection scenarios, ensuring optimal detection performance across a broad spectrum of datasets.
@@ 5.3 Architectural and Training Hyperparameter Importance for F1 Score Optimization
In the SMAP dataset, gaussian_noise with importance 0.53 is paramount, highlighting the model’s need to handle noisy environmental data effectively. Additionally, batch_size with importance 0.07 influences the model’s performance, reflecting the impact of batch processing on training.
@@
The conducted hyperparameter analysis illuminates the tailored performance of TransNAS-TSAD across varied datasets, with each dataset’s specifics dictating the importance of particular hyperparameters. This not only affirms the model’s adaptability but also its capacity for ongoing enhancement, ensuring it remains adept at confronting the evolving complexities of time series data.
@@ 5.4 Pareto Front Analysis for Model Optimization
Building upon the insights from the hyperparameter importance analysis, the Pareto front optimization results for the TransNAS-TSAD model, as depicted in Figure 4, further elucidate the model’s performance. The Pareto front plots for the NAB, MBA, SMAP, and WADI datasets underscore the delicate balance between model complexity, as measured by the number of parameters, and the model’s effectiveness, as quantified by the F1 score.
For the NAB dataset, the Pareto front indicates a dense congregation of model configurations. A significant number of these configurations exhibit a high parameter count without a corresponding increase in the F1 score, suggesting a potential plateau in performance gains relative to complexity. This observation prompts a critical evaluation of model parsimony, encouraging the selection of simpler models that maintain performance while mitigating the risk of overfitting.
@@
These Pareto fronts provide a visual and quantitative tool for identifying optimal model configurations that achieve a balance between performance and complexity. They serve as a decision-making aid for selecting models that not only excel in accuracy but are also efficient and scalable, aligning with the practical demands of anomaly detection in diverse environments.
The findings from the Pareto front analysis are integral to the ongoing development of the TransNAS-TSAD model within the broader scope of our research. They contribute to the understanding of how different model architectures perform across various datasets, offering a pathway to enhancing the model’s adaptability and ensuring its continued efficacy in the face of evolving data challenges.
@@ 6 Discussion: Addressing Challenges and Exploring Future Directions in TransNAS-TSAD
A pivotal challenge addressed by TransNAS-TSAD is the balance of detection accuracy against computational efficiency. The NSGA-II algorithm plays a crucial role in this regard, effectively managing the equilibrium to ensure optimal model performance, without incurring excessive computational demands. This aspect is particularly crucial in real-world applications where resources are finite and efficiency is paramount.
@@
Looking ahead, several promising avenues for enhancement and innovation present themselves. Enhanced real-time data processing capabilities, particularly for applications in environmental monitoring and industrial control systems, represent a significant area for advancement. Techniques from data assimilation and online learning could be effectively integrated into TransNAS-TSAD to address these challenges. Additionally, the development and implementation of dynamic thresholding strategies, such as the Moving Average Thresholding (MAT) and the incorporation of rolling statistics, offer exciting prospects. These augmentative strategies could pave the way for even more nuanced detection capabilities within the TransNAS-TSAD framework.
@@ 7 Conclusion
TransNAS-TSAD emerges as a groundbreaking framework in the realm of time series anomaly detection, effectively integrating transformer architecture with neural architecture search and NSGA-II optimization. The model’s superior performance, evidenced across diverse datasets, demonstrates its robustness in accurately detecting anomalies in both univariate and multivariate time series data.
The strategic application of NSGA-II within TransNAS-TSAD skillfully addresses the challenge of balancing accuracy with computational efficiency, a critical consideration in practical applications. Moreover, the incorporation of advanced adversarial learning paradigms contributes to its ability to detect nuanced anomalies, a notable advancement over traditional methods.
TransNAS-TSAD’s development signifies a notable contribution to the field of anomaly detection, enhancing the capabilities and efficiency of current methodologies. As the landscape of data analysis continues to evolve, the principles and approaches embodied in TransNAS-TSAD will undoubtedly inspire further advancements in machine learning, paving the way for innovative applications in various sectors. The framework not only sets a new benchmark in anomaly detection but also opens up exciting prospects for future research, particularly in enhancing real-time processing and integrating human-centric approaches to machine learning.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program. We would like to extend our gratitude to the providers of the data sets used in this work, whose contributions were invaluable to our research. The authors would like to thank the US Geological Survey (USGS) for offering the domain expertise that was crucial to identify the peak anomaly types that are of practical importance.

### 04 Mar 2024 v3

@@ 5 Results
. The EACS is calculated by normalizing and weighting F1 scores (0.4), training times (0.4), and parameter counts (0.2), demonstrating the balanced efficiency and accuracy of TransNAS-TSAD against benchmark models across multiple datasets,label=Tab: EACS,nofloat=table Method NAB UCR MBA SMAP F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS LSTM-NDT 65.31 23.40 1710 0.8424 52.31 11.14 1765 0.7985 94.56 47.80 18641 0.9520 78.79 36.43 29146 0.8459 DAGMM 74.53 64.50 1266 0.8467 53.37 27.50 1214 0.7919 96.83 92.36 2448 0.9394 88.80 26.60 7266 0.9050 OmniAnomaly 74.42 88.70 13717 0.8250 83.46 38.21 13025 0.8863 92.25 197.66 25474 0.8649 87.28 66.77 16813 0.8230 MSCRED 75.02 503.60 1237377 0.3001 54.41 559.60 128655 0.2176 96.23 774.99 2441778 0.3849 86.64 25.70 8237452 0.6982 MAD-GAN 77.52 53.90 838 0.8671 85.38 31.50 822 0.9177 96.89 229.78 1877 0.8688 86.54 41.26 6718 0.8683 USAD 74.42 65.40 1359 0.8455 89.52 54.23 1359 0.9172 94.43 191.26 1609 0.8789 84.19 49.22 7395 0.8439 CAE-M 79.68 33.60 7229 0.8909 69.81 22.10 7365 0.8520 91.54 89.66 15411 0.9186 88.27 212.51 7229 0.5529 GDN 79.98 131.60 574 0.8153 68.94 64.70 566 0.8286 93.32 203.45 1106 0.8682 85.18 104.50 2974 0.7440 TranAD 93.64 3.37 615 0.9718 94.07 1.14 619 0.9745 97.80 5.11 1298 0.9885 89.15 5.62 62271 0.9445 TransNAS-TSAD 94.11 2.70 285 0.9742 99.10 2.24 1690 0.9922 98.61 2.27 593 0.9932 94.60 3.19 39045 0.9734 Method MSL SWaT WADI SMD F1 % Training Time (sec) # Params EACS F1 Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS F1 % Training Time (sec) # Params EACS LSTM-NDT 77.21 37.80 61856 0.8887 66.70 41.50 72854 0.7957 2.71 422.67 133712 0.5753 90.42 460.70 42853 0.9100 DAGMM 84.82 25.66 18756 0.9258 81.28 29.66 16558 0.8748 14.12 246.32 39387 0.6498 94.91 337.90 10516 0.9420 OmniAnomaly 87.65 29.69 39753 0.9348 81.31 40.20 36541 0.8568 42.60 361.60 39753 0.7794 94.01 311.20 22941 0.9412 MSCRED 93.63 40.70 18476728 0.7535 80.72 236.80 14227264 0.3229 37.41 1884.25 75530 0.6716 84.14 349.77 14237356 0.6978 MAD-GAN 91.69 32.64 17446 0.9497 80.65 34.70 15958 0.8638 35.88 321.20 35764 0.7500 91.50 424.60 9903 0.9188 USAD 88.22 29.60 14859 0.9374 81.43 35.81 12502 0.8651 30.56 389.72 32859 0.7214 94.95 328.77 10609 0.9432 CAE-M 87.33 774.60 204687 0.5471 81.01 71.22 194525 0.8010 41.17 7724.92 388905 0.3503 93.67 3606.90 114219 0.5731 GDN 95.91 121.80 153541 0.9191 81.01 83.55 115442 0.7813 42.60 6047.12 287082 0.4559 83.42 1000.40 7255 0.8226 TranAD 94.94 6.88 272181 0.9733 81.53 2.10 31336 0.9221 49.51 177.60 1378173 0.6645 96.05 56.70 135110 0.9760 TransNAS-TSAD 94.82 5.44 236491 0.9739 83.14 1.60 13464 0.9297 84.35 227.50 1701515 0.7883 99.81 52.40 132050 0.9916
in time series anomaly detection.,label=tab:Best model params,nofloat=table Arch and Training Params SMAP UCR MBA SWaT MSL SMD NAB WADI Learning Rate 0.0002128 0.0019997 0.003441 0.0000403 0.006192 0.0002273 0.006924 0.0015406 Dropout Rate 0.2353 0.4474 0.1795 0.3836 0.3730 0.1554 0.4560 0.2886 Dim Feedforward 101 124 41 42 121 86 10 24 Batch Size 32 48 16 32 32 48 96 128 Encoder Layers 2 1 3 2 2 3 1 2 Decoder Layers 1 2 1 3 3 1 2 1 Activation Func. sigmoid leaky_relu tanh tanh leaky_relu sigmoid relu leaky_relu Time Warping False True True True True False False True Time Masking True True False True True False True False Gaussian Noise 0.000151 0.027956 0.007925 0.058019 0.000428 0.000628 0.000119 0.001042 Linear Embedding False True False False True True True True Phase Type 2phase iterative iterative 2phase 2phase 2phase iterative 2Phase Self Conditioning False True False True True False True False Layer Norm False False False True True True False False Pos. Enc. Type sinusoidal fourier fourier sinusoidal fourier sinusoidal sinusoidal sinusoidal FFN Layers 1 1 1 3 2 3 1 1 Attn Heads 25 1 2 51 55 38 1 127 Window Size 10 20 14 22 26 12 18 26
5.1 Optimal Model Configurations Identified by TransNAS-TSAD
@@
Table LABEL:tab:Best_model_params showcases the culmination of TransNAS-TSAD’s NSGA-II optimization process, highlighting the training and architectural hyperparameters of models that achieved the best balance between high F1 scores and low parameter counts. These configurations represent models uniquely suited to each dataset’s anomaly detection needs.
@@
A key observation is the diversity in architectural parameters across datasets, reflecting TransNAS-TSAD’s adeptness in tailoring models to specific data characteristics. For example, in the SMAP dataset, the model employs a relatively simple architecture with fewer encoder and decoder layers, which is effective in handling the environmental time series data of this dataset. In contrast, the WADI dataset, known for its complex sensor network, necessitates a more intricate model structure, evident in its higher number of attention heads and the use of a two-phase reconstruction and refinement strategy.
@@
These results from TransNAS-TSAD’s NAS process validate the framework’s ability to not only fine-tune hyperparameters for operational efficiency, but also to intricately adapt its architectural design to the nuanced demands of various time series anomaly detection scenarios, ensuring optimal detection performance across a broad spectrum of datasets.
@@ 5.2 Architectural and Training Hyperparameter Importance for F1 Score Optimization
In the SMAP dataset, gaussian_noise with importance 0.53 highlights the model’s ability to handle noisy environmental data effectively. Additionally, batch_size with importance 0.07 influences the model’s performance, reflecting the impact of batch processing on training.
@@
This hyperparameter analysis illuminates the tailored performance of TransNAS-TSAD across varied datasets, with each dataset’s specifics dictating the importance of particular hyperparameters. This not only affirms the model’s adaptability but also its capacity for ongoing enhancement, ensuring it remains adept at confronting the evolving complexities of time series data.
@@ 5.3 Pareto Front Analysis for Model Optimization
Building upon the insights from the hyperparameter importance analysis, the Pareto front optimization results for the TransNAS-TSAD model are depicted in Figure 4. The Pareto front plots for the NAB, MBA, SMAP, and WADI datasets underscore the delicate balance between model complexity, as measured by the number of parameters, and the model’s effectiveness, as quantified by the F1 score.
For the NAB dataset, the Pareto front indicates a dense congregation of model configurations. A significant number of these configurations exhibit a high parameter count without a corresponding increase in the F1 score, suggesting a potential plateau in performance gains relative to complexity. This observation prompts a critical evaluation of model parsimony, encouraging the selection of simpler models that maintain performance while mitigating the risk of over-fitting.
@@
These Pareto fronts provide a visual and quantitative tool for identifying optimal model configurations that achieve a balance between performance and complexity. They serve as a decision-making aid for selecting models that align with the practical demands of anomaly detection in diverse environments.
The findings from the Pareto analysis are integral to the ongoing development of the TransNAS-TSAD model within the broader scope of our research. They contribute to understanding how different model architectures perform across various datasets, and offer a pathway to enhancing the model’s adaptability and ensuring its continued efficacy in the face of evolving data challenges.
@@ 6 Discussion: Addressing Challenges and Exploring Future Directions in TransNAS-TSAD
A pivotal challenge addressed by TransNAS-TSAD is the balance between detection accuracy and computational efficiency. The NSGA-II algorithm plays a critical role in optimizing model performance, without incurring excessive computational demands. This aspect is particularly crucial in real-world applications where resources are finite and efficiency is paramount.
@@
Looking ahead, several promising avenues for enhancement and innovation present themselves. Enhanced real-time data processing capabilities, particularly for applications in environmental monitoring and industrial control systems, represent a significant area for advancement. Techniques from data assimilation and online learning could be effectively integrated into TransNAS-TSAD to address these challenges. Additionally, the development and implementation of dynamic thresholding strategies, such as the Moving Average Thresholding (MAT) and the incorporation of rolling statistics, offer exciting prospects.
@@ 7 Conclusion
TransNAS-TSAD signifies a notable contribution in the realm of time series anomaly detection by merging transformer architecture with neural architecture search and NSGA-II optimization, achieving superior performance across diverse univariate and multivariate datasets. Its robustness in accurately detecting anomalies highlights the effectiveness of this integration, effectively addressing the dual challenge of maintaining accuracy while ensuring computational efficiency—key for practical applications. This framework is further distinguished by its incorporation of advanced adversarial learning paradigms, enabling the detection of nuanced anomalies and marking a significant step forward from traditional methods. As the landscape of data analysis evolves, TransNAS-TSAD not only sets new performance benchmarks in anomaly detection but also opens up exciting avenues for future research, particularly in enhancing real-time processing and integrating human-centric approaches to machine learning. The principles and approaches embodied in TransNAS-TSAD are paving the way for innovative applications across various sectors, shaping the future of machine learning with its groundbreaking methodologies.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program. We would like to extend our gratitude to the providers of the data sets used in this work, whose contributions were invaluable to our research.